10. Find the coefficient of x20 in A(x)=(x10+x11+.....+x25)(x1+x2+....+x15)5(x20+x21+.....+x45)


1
Expert's answer
2022-02-25T07:56:13-0500

Given, "A(x)=(x^{10}+x^{11}+\cdots+x^{25})(x^1+x^2+\cdots+x^{15})^{5}(x^{20}+x^{21}+\cdots+x^{45})"

"\begin{aligned}
\text{Coefficient of $x^{20}$~in}~A(x) &= \text{Coefficient of $x^{20}$~in~} (x^{10}+x^{11}+\cdots+x^{25})\\ &\qquad(x^1+x^2+\cdots+x^{15})^{5}(x^{20}+x^{21}+\cdots+x^{45})\\
&= \text{Coefficient of $x^{20}$~in~} [x^{10} (1+x+\cdots+x^{15})\cdot \\&\qquad x^{5}(1+x+\cdots+x^{14})^{5}\cdot x^{20}(1+x+\cdots+x^{25})]\\
&=\text{Coefficient of $x^{20}$~in~} x^{35} [(1+x+\cdots+x^{15})\cdot \\&\qquad (1+x+\cdots+x^{14})^{5}\cdot(1+x+\cdots+x^{25})]\\
\end{aligned}"

The first term in the expansion of"x^{35} [(1+x+\cdots+x^{15})\cdot (1+x+\cdots+x^{14})^{5}\cdot(1+x+\cdots+x^{25})]" will be "x^{35}" and hence coefficient of "x^{20}" in A(x) = 0, since negative powers cannot exist in the expansion of "(1+x+\cdots+x^{15})\cdot (1+x+\cdots+x^{14})^{5}\cdot(1+x+\cdots+x^{25})."